OpenStudy (anonymous):

Let f be the function defined by f(x)= -2 + lnx^2. a. for what real numbers x is f defined? b. find the zeros of f. c.write an equation for the line tangent to the graph of f at x=1

OpenStudy (anonymous):

defined for all numbers except \(0\) because you cannot take the log of zero

OpenStudy (anonymous):

\[-2+\ln(x^2)=0\] \[\ln(x^2)=2\] \[2\ln(x)=2\] \[\ln(x)=1\]\[x=e\]